OpenStudy (anonymous):

Solve this system of equations using substitution. Put your answer in ordered pair form, (x,y) a=3b-4 a+b=16

OpenStudy (anonymous):

a-3b=-4 a+b=16 subtract the equation -4b=-20 b=5 a+b=16 a+5 = 16 a= 11 (11,5)

OpenStudy (anonymous):

@Kittyrae15 understood

OpenStudy (anonymous):

Yes, thanks again:p